# HiCL: Hippocampal-Inspired Continual Learning

### v2

@@ Abstract
We propose HiCL, a novel hippocampal-inspired dual-memory continual learning architecture designed to mitigate catastrophic forgetting by using elements inspired by the hippocampal circuitry. Our system encodes inputs through a grid-cell-like layer, followed by sparse pattern separation using a dentate gyrus-inspired module with top-$k$ sparsity. Episodic memory traces are maintained in a CA3-like autoassociative memory. Task-specific processing is dynamically managed via a DG-gated mixture-of-experts mechanism, wherein inputs are routed to experts based on cosine similarity between their normalized sparse DG representations and learned task-specific DG prototypes computed through online exponential moving averages. This biologically grounded yet mathematically principled gating strategy enables differentiable, scalable task-routing without relying on a separate gating network, and enhances the model’s adaptability and efficiency in learning multiple sequential tasks. Cortical outputs are consolidated using Elastic Weight Consolidation weighted by inter-task similarity. Crucially, we incorporate prioritized replay of stored patterns to reinforce essential past experiences. Evaluations on standard continual learning benchmarks demonstrate the effectiveness of our architecture in reducing task interference, achieving near state-of-the-art results in continual learning tasks at lower computational costs. Our code is available here https://github.com/kushalk173-sc/HiCL.
@@ Introduction
In continual learning, every new task is like a distinct episode in our life’s unfolding story. Just as the hippocampus rapidly encodes unique experiences into episodic memory, tagging them with contextual and temporal cues, a continual learner must capture each task’s defining features before they fade. The brain then replays these episodes—during sleep or quiet wakefulness—to gradually integrate them into neocortical networks, preventing older memories from being overwritten by new ones. We leverage this complementary learning systems principle by treating each task as an episodic trace: we store compact representations in a replay buffer and periodically revisit them during training, mirroring how hippocampal replay consolidates previous and precious episodes into lasting knowledge (McClelland, McNaughton, and O’Reilly 1995).
Figure 1: Schematic of the HiCl architecture overlaid with the hippocampal trisynaptic circuit. The entorhinal cortex (layers II/III) projects to the dentate gyrus (DG), which performs pattern separation. DG outputs feed into CA3, an autoassociative network responsible for pattern completion. CA3 then projects to CA1, which integrates inputs from both CA3 and the entorhinal cortex before sending output to other brain regions. This circuit underlies memory encoding and retrieval processes. The general background here of the hippocampal circuit is based on (Petrantonakis and Poirazi 2014) and adapted from Moser, Kropff, and Moser (2008); Leutgeb et al. (2007); Rolls (2013b); Otmakhova, Otmakhov, and Griffith (2018).
We draw on three core hippocampal subregions—the dentate gyrus (DG), CA3, and CA1—to inspire a DG‑Gated Mixture‑of‑Experts (MoE) architecture, where each subcircuit is instantiated as an expert and inputs are dynamically routed based on their DG activations. This DG‑based gating, eliminates the need for task labels at inference and directly tackles catastrophic forgetting.
@@
  • We achieved competitive accuracy at a fraction of the computational cost. On benchmarks like Split CIFAR-10, our LeNet-based model demonstrates that a neuroscience-aligned architecture can be both highly effective and computationally efficient in a continual learning setting.
@@
The CLS theory posits a dual-memory system: a fast hippocampal module for episodic learning and a slow neocortical system for generalization (McClelland, McNaughton, and O’Reilly 1995; Kemker et al. 2018). This foundational concept has influenced many biologically inspired CL frameworks. Examples include hybrid cortico-hippocampal networks that combine feedforward and spiking layers (Rothschild, Eban, and Frank 2017), DualNet’s distinct fast/slow learners (Pham, Liu, and Hoi 2021), FearNet’s use of generative replay (Kemker and Kanan 2017), and Triple Memory Networks that model hippocampus-PFC interactions with replay and consolidation (Wang et al. 2021).
@@
Many CL frameworks are categorized as parameter isolation, rehearsal based, or regularization based (De Lange et al. 2021). Regularization methods add a loss penalty to prevent changes to weights important for past tasks, as seen in Elastic Weight Consolidation (EWC) (Kirkpatrick et al. 2017), Learning without Forgetting (LwF) (Li and Hoiem 2017), and Synaptic Intelligence (SI) (Zenke, Poole, and Ganguli 2017). Parameter isolation methods assign distinct weights to different tasks, using techniques such as network expansion in Progressive Nets (Rusu et al. 2016) or hard attention masks (Serrà et al. 2018; Madotto et al. 2021; Wang et al. 2024). Rehearsal methods use a buffer of past data (Rebuffi et al. 2017; Riemer et al. 2018; Zhang, Wang, and Yang 2022) or a generative model (Shin et al. 2017; Su et al. 2019; Skierś and Deja 2025). Hybrid models such as Progress & Compress (Schwarz et al. 2018) and CLEAR (Riemer et al. 2018) combine these strategies effectively, and our approach synergizes these ideas by using architectural modularity, sparsity, and replay.
@@
MoE models enable modular specialization by routing inputs to specialized subnetworks (Yuksel, Wilson, and Gader 2012; Shazeer et al. 2017), a principle recently applied to CL (Li et al. 2025; Le et al. 2024) and large-scale models (Guo et al. 2025; Cai et al. 2025). Stability is often improved via adaptive routing, for instance by annealing a routing temperature (Nie et al. 2022; Fedus, Zoph, and Shazeer 2022). Our DG-Gated MoE is distinct in that expert selection emerges via cosine similarity to prototypes, bypassing an explicit learned gating network.
@@
Storing class prototypes instead of raw data is a memory-efficient rehearsal strategy pioneered by iCaRL (Rebuffi et al. 2017). Subsequent work has combined prototype anchoring with contrastive objectives to preserve the geometry of the feature space across tasks (Caccia et al. 2021; Aghasanli, Li, and Angelov 2025). We follow this line, maintaining online exponential moving average (EMA) prototypes and applying a push-pull contrastive loss during replay.
@@
Our design is grounded in the hippocampal trisynaptic circuit (DG $\rightarrow$ CA3 $\rightarrow$ CA1) and the theory of complementary learning systems (CLS) (McClelland, McNaughton, and O’Reilly 1995; Kemker et al. 2018). We operationalize these motifs as inductive scaffolds: structured encoding, sparse separation, nonlinear refinement/completion, and integration, which together feed into prototype-based routing and consolidation. Replay-inspired mechanisms and prototype anchoring provide retention and interference mitigation, while modular gating arises naturally from the quasi-orthogonal codes produced by DG-like separation. Learning is organized as a two-phase dual-memory schedule—specialization followed by consolidation—whose detailed losses are defined in the Architecture section.
@@
$\mathbf{g}_{m}=\sin(\mathbf{W}_{m}\,\mathbf{f}+\phi_{m}),\quad m=1,\dots,4,\qquad\mathbf{g}=[\mathbf{g}_{1};\dots;\mathbf{g}_{4}].$
@@
This top-$k$ mechanism approximates inhibitory microcircuitry to produce quasi-orthogonal sparse codes, creating natural modularization and enabling routing without explicit task labels, akin to capacity partitioning in MoE and gating literature (Yuksel, Wilson, and Gader 2012; Shazeer et al. 2017).
@@
$\mathbf{p}_{\mathrm{comp}}=\mathrm{LayerNorm}\bigl(\mathrm{ReLU}(\mathbf{W}_{2}\,\mathrm{ReLU}(\mathbf{W}_{1}\,\mathbf{p}_{\mathrm{sep}}+\mathbf{b}_{1})+\mathbf{b}_{2})\bigr),$
@@ Results
Size Task IL ↑ Class IL ↑ FLOPs (M) ↓ Task IL ↑ Class IL ↑ FLOPs (M) ↓
@@
Table 1: Comparison of accuracy (%), and FLOPs (M) on Split CIFAR-10 and Split TinyImageNet under different memory and compute settings. The baselines have been updated based on new results for Split CIFAR-10.
@@
  • Split CIFAR-10 (Krizhevsky, Hinton et al. 2009): The 10 classes are divided into 5 sequential tasks, with 2 new classes introduced in each task.
@@
Figure 3: DG-Gated Routing is Enabled by Effective Feature Separation. (a) The Class-IL routing matrix shows that inputs are correctly routed to their corresponding expert. For example, inputs from classes 0 and 1 (Task 0) are primarily routed to Expert 0. (b) This routing is enabled by strong pattern separation in the DG layer, where the Jaccard similarity between patterns from different tasks (’Inter-Task’) is significantly lower than within the same task (’Intra-Task’).
@@ Discussion
(a) CIFAR-10 t-SNE
(b) CIFAR-10 Prototype Similarity
(c) Tiny-ImageNet t-SNE
(d) Tiny-ImageNet Prototype Similarity
Figure 4: DG Feature Separation and Prototype Orthogonality. t-SNE plots (a, c) and similarity matrices (b, d) for Split CIFAR-10 and Split Tiny-ImageNet.
This separation is quantified by the prototype similarity matrices in Figure 4b and 4d. The strong diagonal shows that each prototype is similar to itself, while the low off-diagonal values (all close to zero) confirm that the prototypes for different tasks are nearly orthogonal. Therefore, the routing decision, a simple cosine similarity check against the prototypes is not just a superficial pattern match. It is a measure of representational resonance. The input is naturally routed to the expert whose internal feature vocabulary is best suited to understand and process it, a choice made simple and reliable by the clear, quantifiable separation.
@@
As shown in Figure 5, HiCL degrades gracefully as the replay buffer size shrinks. For instance, reducing the buffer from 500 to 100 examples per task causes Task-IL accuracy to drop by only 2.7 percentage points (from 92.8% to 90.1%). This robustness stems from the DG’s pattern separation: each stored exemplar occupies a distinct sparse subspace, maximizing its utility during replay. Moreover, the model’s feature-distillation and EWC regularization further stabilize representations, mitigating the effect of limited memory. These results suggest that efficient episodic encoding—rather than large buffer capacity—is a key driver of continual-learning performance.
Figure 5: Accuracy vs. Memory Size on Split CIFAR-10. The plot shows the final Task-IL Accuracy (top band), Class-IL Accuracy (blue line), and the model’s internal routing accuracy (orange dashed line) as the number of stored examples per task (memory size) varies. HiCL maintains high Task-IL accuracy and shows graceful degradation in Class-IL accuracy even with smaller buffer sizes.

### v3

@@ Abstract
We propose HiCL, a novel hippocampal-inspired dual-memory continual learning architecture designed to mitigate catastrophic forgetting by using elements inspired by the hippocampal circuitry. Our system encodes inputs through a grid-cell-like layer, followed by sparse pattern separation using a dentate gyrus-inspired module with top-$k$ sparsity. Episodic memory traces are maintained sin a CA3-like autoassociative memory. Task-specific processing is dynamically managed via a DG-gated mixture-of-experts mechanism, wherein inputs are routed to experts based on cosine similarity between their normalized sparse DG representations and learned task-specific DG prototypes computed through online exponential moving averages. This biologically grounded yet mathematically principled gating strategy enables differentiable, scalable task-routing without relying on a separate gating network, and enhances the model’s adaptability and efficiency in learning multiple sequential tasks. Cortical outputs are consolidated using Elastic Weight Consolidation weighted by inter-task similarity. Crucially, we incorporate prioritized replay of stored patterns to reinforce essential past experiences. Evaluations on standard continual learning benchmarks demonstrate the effectiveness of our architecture in reducing task interference, achieving near state-of-the-art results in continual learning tasks at lower computational costs.
Codehttps://github.com/kushalk173-sc/HiCL
Extended versionhttps://arxiv.org/pdf/2508.16651
@@ Introduction
In continual learning, every new task is like a distinct episode in our life’s unfolding story. Just as the hippocampus rapidly encodes unique experiences into episodic memory, tagging them with contextual and temporal cues, a continual learner must capture each task’s defining features before they fade. The brain then replays these episodes—during sleep or quiet wakefulness—to gradually integrate them into neocortical networks, preventing older memories from being overwritten by new ones. We leverage this complementary learning systems principle by treating each task as an episodic trace: we store compact representations in a replay buffer and periodically revisit them during training, mirroring how hippocampal replay consolidates previous and precious episodes into lasting knowledge (McClelland et al. 1995).
Figure 1: Schematic of the HiCl architecture overlaid with the hippocampal trisynaptic circuit. The entorhinal cortex (layers II/III) projects to the dentate gyrus (DG), which performs pattern separation. DG outputs feed into CA3, an autoassociative network responsible for pattern completion. CA3 then projects to CA1, which integrates inputs from both CA3 and the entorhinal cortex before sending output to other brain regions. This circuit underlies memory encoding and retrieval processes. The general background here of the hippocampal circuit is based on (Petrantonakis and Poirazi 2014) and adapted from Moser et al. (2008); Leutgeb et al. (2007); Rolls (2013b); Otmakhova et al. (2018).
We draw on three core hippocampal subregions—the dentate gyrus (DG), CA3, and CA1—to inspire a DG‑Gated Mixture‑of‑Experts (MoE) architecture(as shown in Figure 1, where each subcircuit is instantiated as an expert and inputs are dynamically routed based on their DG activations. This DG‑based gating, eliminates the need for task labels at inference and directly tackles catastrophic forgetting.
@@
  • We achieved competitive accuracy at a fraction of the computational cost. On benchmarks like Split CIFAR-10, our model demonstrates that a neuroscience-aligned architecture can be both computationally effective and efficient in a continual learning setting.
@@
The CLS theory posits a dual-memory system: a fast hippocampal module for episodic learning and a slow neocortical system for generalization (McClelland et al. 1995; Kemker et al. 2018). This foundational concept has influenced many biologically inspired CL frameworks. Examples include hybrid cortico-hippocampal networks that combine feedforward and spiking layers (Rothschild et al. 2017), DualNet’s distinct fast/slow learners (Pham et al. 2021), FearNet’s use of generative replay (Kemker and Kanan 2017), and Triple Memory Networks that model hippocampus-PFC interactions with replay and consolidation (Wang et al. 2021).
@@
Many CL frameworks are categorized as parameter isolation, rehearsal based, or regularization based (De Lange et al. 2021). Regularization methods add a loss penalty to prevent changes to weights important for past tasks, as seen in Elastic Weight Consolidation (EWC) (Kirkpatrick et al. 2017), Learning without Forgetting (LwF) (Li and Hoiem 2017), and Synaptic Intelligence (SI) (Zenke et al. 2017). Parameter isolation methods assign distinct weights to different tasks, using techniques such as network expansion in Progressive Nets (Rusu et al. 2016) or hard attention masks (Serrà et al. 2018; Madotto et al. 2021; Wang et al. 2024). Rehearsal methods use a buffer of past data (Rebuffi et al. 2017; Riemer et al. 2018; Zhang et al. 2022) or a generative model (Shin et al. 2017; Su et al. 2019; Skierś and Deja 2025). Hybrid models such as Progress & Compress (Schwarz et al. 2018) and CLEAR (Riemer et al. 2018) combine these strategies effectively, and our approach synergizes these ideas by using architectural modularity, sparsity, and replay.
@@
MoE models enable modular specialization by routing inputs to specialized subnetworks (Yuksel et al. 2012; Shazeer et al. 2017), a principle recently applied to CL (Li et al. 2025; Le et al. 2024) and large-scale models (Guo et al. 2025; Cai et al. 2025). Stability is often improved via adaptive routing, for instance by annealing a routing temperature (Nie and others 2022; Fedus et al. 2022). Our DG-Gated MoE is distinct in that expert selection emerges via cosine similarity to prototypes, bypassing an explicit learned gating network.
@@
Storing class prototypes instead of raw data is a memory-efficient rehearsal strategy pioneered by iCaRL (Rebuffi et al. 2017). Subsequent work has combined prototype anchoring with contrastive objectives to preserve the geometry of the feature space across tasks (Caccia et al. 2021; Aghasanli et al. 2025). We follow this line, maintaining online exponential moving average (EMA) prototypes and applying a push-pull contrastive loss during replay.
@@
Our design is grounded in the hippocampal trisynaptic circuit (DG $\rightarrow$ CA3 $\rightarrow$ CA1) and the theory of complementary learning systems (CLS) (McClelland et al. 1995; Kemker et al. 2018). We operationalize these motifs as inductive scaffolds: structured encoding, sparse separation, nonlinear refinement/completion, and integration, which together feed into prototype-based routing and consolidation. Replay-inspired mechanisms and prototype anchoring provide retention and interference mitigation, while modular gating arises naturally from the quasi-orthogonal codes produced by DG-like separation. Learning is organized as a two-phase dual-memory schedule—specialization followed by consolidation—whose detailed losses are defined in the Architecture section.
@@
$\mathbf{g}_{m}$ $=\sin(\mathbf{W}_{m}\,\mathbf{f}+\phi_{m}),\quad m=1,\dots,4,\mathbf{g}=[\mathbf{g}_{1};\dots;\mathbf{g}_{4}].$
@@
This top-$k$ mechanism approximates inhibitory microcircuitry to produce quasi-orthogonal sparse codes, creating natural modularization and enabling routing without explicit task labels, akin to capacity partitioning in MoE and gating literature (Yuksel et al. 2012; Shazeer et al. 2017).
@@
$\mathbf{p}_{\mathrm{comp}}$ $=\mathrm{LN}\bigl(\mathrm{ReLU}(\mathbf{W}_{2}\,\mathrm{ReLU}(\mathbf{W}_{1}\,\mathbf{p}_{\mathrm{sep}}$
$\quad+\mathbf{b}_{1})+\mathbf{b}_{2})\bigr)$
@@ Results
Size Task IL ↑ Class IL ↑ FLOPs ↓ Task IL ↑ Class IL ↑ FLOPs↓
@@
Table 1: Comparison of accuracy (%), and FLOPs (M) on Split CIFAR-10 and Split TinyImageNet under different memory and compute settings. The baselines have been updated based on new results for Split CIFAR-10. FLOPs are reported in millions.
@@
  • Split CIFAR-10 (Krizhevsky et al. 2009): The 10 classes are divided into 5 sequential tasks, with 2 new classes introduced in each task.
@@
Figure 3: DG-Gated Routing is Enabled by Effective Feature Separation. The Class-IL routing matrix shows that inputs are correctly routed to their corresponding expert. For example, inputs from classes 0 and 1 (Task 0) are primarily routed to Expert 0. This routing is enabled by strong pattern separation in the DG layer, where the Jaccard similarity between patterns from different tasks (’Inter-Task’) is significantly lower than within the same task (’Intra-Task’).
@@ Discussion
CIFAR-10 t-SNE
Figure 4: DG Feature Separation and Prototype Orthogonality. t-SNE plot for Split CIFAR-10.
Tiny-ImageNet t-SNE
Figure 5: DG Feature Separation and Prototype Orthogonality. t-SNE plot for Split Tiny-ImageNet.
This separation is illustrated by the two dimensional t-SNE embeddings in Figures 4 and 5. Each colored group corresponds to a learned prototype, and the points form compact clusters with clear gaps between groups. The minimal overlap across clusters indicates that prototypes associated with different tasks occupy distinct regions of the representation space, which is consistent with low cross task similarity and reduced interference during continual learning. Therefore, the routing decision, a simple cosine similarity check against the prototypes is not just a superficial pattern match. It is a measure of representational resonance. The input is naturally routed to the expert whose internal feature vocabulary is best suited to understand and process it, a choice made simple and reliable by the clear, quantifiable separation.
@@
As shown in Table 2, HiCL degrades gracefully as the replay buffer size shrinks. For instance, reducing the buffer from 500 to 100 examples per task causes Task-IL accuracy to drop by only 2.7 percentage points (from 92.8% to 90.1%). This robustness stems from the DG’s pattern separation: each stored exemplar occupies a distinct sparse subspace, maximizing its utility during replay. Moreover, the model’s feature-distillation and EWC regularization further stabilize representations, mitigating the effect of limited memory. These results suggest that efficient episodic encoding—rather than large buffer capacity—is a key driver of continual-learning performance.
Memory Size Task-IL Class-IL Routing
100 90.1 53.4 56.0
200 92.6 60.6 63.2
300 91.5 63.1 65.6
400 92.2 64.2 67.7
500 92.8 65.4 67.8
600 91.3 64.6 67.6
1000 92.5 67.2 69.1
5000 92.8 63.7 64.1
Table 2: Accuracy vs Memory Size (Task-IL, Class-IL, Routing)